OpenStudy (anonymous):

3(x − 1) − (x + 5) = 2

OpenStudy (anonymous):

Have you heard of Distributive Property ??

OpenStudy (gw2011):

The equation simplifies to: 3x-3-x-5+2 2x-8=2 2x=10 x=5

OpenStudy (anonymous):

@waterineyes yes i have i just dont do well im math

OpenStudy (anonymous):

@gw2011 thanks

OpenStudy (gw2011):

You're welcome.

OpenStudy (anonymous):

If you have heard then try to implement it on the questions.. You will surely do well if you will practice..

OpenStudy (anonymous):

Distributive property says: \[a(b+c) \implies ab + ac\]

OpenStudy (anonymous):

Now : \[3(x-1) = 3 \times x - 3 \times 1 \implies 3x - 3\] Here basically we distribute 3 equally to x and -1..

OpenStudy (anonymous):

then its 3x-3-x-5+2

OpenStudy (anonymous):

Actually there is not + 2 in between there will equal to sign that is = 2: 3x - 3 -x - 5 = 2 @gw2011 has done a typing mistake there..

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

well thanks for the help

OpenStudy (anonymous):

but his next steps are good and fine..

OpenStudy (anonymous):

ok

OpenStudy (gw2011):

Sorry, when I wrote 3x-3-x-5+2, it should have been 3x-3-x-5=2. I hit the wrong key originally.

OpenStudy (anonymous):

It is okay @gw2011